OpenStudy (anonymous):

Jenna buys 12 lb of meat. Some of the meat is chicken and costs $1.90/lb. The rest is beef, which costs $4.30/lb. She spends a total of $34.80. Which equation models this situation? Let c represent the number of pounds of chicken she buys. A. (1.9 + 4.3)(12 – c) = 34.8 B. 4.3c + 1.9(c + 14) = 34.8 C. 4.3c + 1.9(12) = 34.8 D. 1.9c + 4.3(12 − c) = 34.8

OpenStudy (displayerror):

We're told that chicken costs $1.90/lb. How much would 2 lbs of chicken cost? Now, how much would c pounds of chicken cost?

OpenStudy (anonymous):

1.9c + 4.3(12 - c) = 34.8

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

so D

OpenStudy (displayerror):

Yep! If we have two pounds of chicken, the cost would be \[\frac{$1.9}{\text{pound of chicken}} \times \text{2 pounds of chicken} = $1.9 \times 2 = $3.8\] So if we have "c" pounds of chicken, the cost would be \[$1.9 \times \text{c}\] Then the amount of meat that is not chicken (out of the total 12 pounds of meat) would be \[12 - c \ \ \text{pounds of beef}\] And the cost of beef, following what we did for the cost of chicken, would be \[\frac{$4.3}{\text{pound of beef}} \times \left( 12 - c \ \text{pounds of beef}\right)\]

OpenStudy (theedwardsfamily):

@amysparkly12 the answer's D